OpenStudy (anonymous):

-3x^2+6x-2 Express the quadratic function in standard form

OpenStudy (zzr0ck3r):

What is standard form?

OpenStudy (anonymous):

(x-h)^2+k

OpenStudy (anonymous):

convert this equation in square form.

OpenStudy (anonymous):

You need to get a perfect square. The formula for a perfect square is (a+b)^2=a^2+b^2+2ab. Try summing and subtracting the same thing to get something like this.

OpenStudy (anonymous):

im confused...

OpenStudy (zzr0ck3r):

It is hard for me to explain how to complete the square on this site but I will try. we have -3x^2+6x-2 what we need to do is get it to the form like you showed. I like to set them equal to 0, because it helps me think about the factors, and it works with my algorithm. you still with me @brent0423 ?

OpenStudy (anonymous):

-3x^2+6x-2 factor out the -3 from the first two terms: -3x^2+6x-2 = -3(x^2 -2x ) - 2 to make what's in the parenthesis a perfect square trinomial, add 1. but in doing so, you've actually subtracted 3 because of the factor of -3 outside so you need to keep the expression the same by ADDing 3 -3x^2+6x-2 = -3(x^2 -2x + 1 ) - 2 + 3 = -3(x - 1)^2 - 2 + 3 = -3(x - 1)^2 + 1